OpenStudy (anonymous):

Jill arranges quarters in stacks of increasing powers of 2, starting with the zeroth power (20). If she ends up with six stacks, how many coins has she used?

OpenStudy (anonymous):

Stack 1: 2^0 = 1 quarter Stack 2: 2^1 = 2 quarters Keep going like this until Stack 6 = 2^5 = 32 quarters Add up 1 + 2 + 4 + 8 + 16 + 32 = 63 quarters.

OpenStudy (anonymous):

@Klip when you go to pet a dog.. . . . . . . . . .. . . . . . . . . .. ...... and it growls at you.

OpenStudy (anonymous):

LOL. Right?

OpenStudy (anonymous):

cx